OpenStudy (anonymous):

Find the remainder of the division problem shown below. (x^4+x^3-3x-3)÷(x-2)

OpenStudy (amistre64):

all it takes is for yoo to do longhand division; do they still teach that in schools?

OpenStudy (anonymous):

replace x by 2 and that will be your remainder. you do not have to divide

OpenStudy (amistre64):

you dont HAVE to; but you can :)

OpenStudy (amistre64):

synthetic pops out a good show too

OpenStudy (anonymous):

\[2^4+2^3-3\times 2-3=15\]

OpenStudy (anonymous):

@amistre synthetic is same as evaluating polynomial

OpenStudy (anonymous):

(x^4+x^3-3x-3)÷(x-2) x-2=0 x=2 f(x)=x^4+x^3-3x-3 f(2)=2^4+2^3-3(2)-3 f(2)=16+8-6-3=15

OpenStudy (anonymous):

so for next time i just plug in what ever number is in the problem that is doing the dividing?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

1 1 0 -3 -3 0 2 6 12 18 ----------- 2| 1 3 6 9 15 <-- remainder of 15/(x-2)

OpenStudy (anonymous):

would the answer change if it was (x+2)?